# Discovering Interpretable Programmatic Policies via Multimodal LLM-assisted Evolutionary Search

### v1

@@ 4.3 Generalization via Policy Ensemble
Table 2: Effect of Policy Ensemble on LES Performance
As illustrated in Fig. 5, MLES generates a diverse population of high-performing policies in a single run. This opens up new avenues for improving generalization, particularly by applying ensemble decision-making techniques to these high-performing policies. Table 2 presents the results of a simple ensemble strategy, where discrete actions are voted on and continuous actions are averaged. Here, MLES and EoH denote the versions that employ ensemble decisions based on the entire population of policies.
Despite its simplicity, this ensemble approach significantly enhances test performance for both MLES and EoH, highlighting the potential of ensemble decision-making to boost generalization. Notably, a slight decline in training performance is observed after ensembling. This outcome aligns with the well-known bias-variance tradeoff in machine learning, where ensembling generally improves generalization at the cost of slightly increased bias. In future work, we plan to explore more advanced techniques for enhancing generalization, such as incorporating instance feature awareness and adaptive collaborative decision-making. Overall, MLES shows inherent advantages and substantial potential in addressing the generalization issues.
A statistical analysis of Table 1 and an investigation into how different forms of IBE and prompts influence the policy discovery process are presented in Appendix D.
5 Conclusion
This paper presents Multimodal Large Language Model-assisted Evolutionary Search (MLES), which combines multimodal large language models with evolutionary computation to efficiently design interpretable programmatic policy for control tasks. By introducing visual feedback-driven behavior analysis, MLES mimics the process of human policy discovery. The framework is demonstrated on two standard control tasks, showcasing its policy discovery capability and efficiency comparable to Proximal Policy Optimization (PPO). Notably, MLES yields transparent control logic with traceable policy design processes. MLES offers a flexible and automated approach that reduces human effort and facilitates knowledge reuse, presenting a new paradigm for the next generation of interpretable control policy discovery.
Appendix A Discussion and Limitation
A.1 Outlook of MLES for Policy Discovery
This paper provides a preliminary demonstration of the Multimodal Large Language Model-assisted Evolutionary Search (MLES) framework’s ability to autonomously discover programmatic policies. As detailed in Appendix F, MLES successfully constructed a data processing pipeline for image data in the Car Racing task, guiding control decisions for steering, throttle, and braking with flawless performance across various tracks. These achievements highlight the promising potential of MLES for future advancements.
Modular feature processing and decision-making
As observed in Appendix F, the programmatic policy discovered by MLES is modular, comprising two main components: feature processing and decision-making. MLES currently utilizes libraries like OpenCV for image data processing, highlighting its ability to autonomously employ such tools for feature processing. This also demonstrates the promising potential for integrating more advanced feature processing modules, such as convolutional neural networks, to further enhance data handling. These advanced modules would then feed processed features into an interpretable control logic for decision-making. Unlike deep reinforcement learning (DRL), where feature processing and decision-making are both black-box and tightly coupled, MLES ensures that even if feature processing is a black box, the decision-making remains interpretable. This clear modularization enables MLES to tackle complex scenarios while maintaining fully verifiable and explainable control logic. Such capabilities are essential for tasks like embodied intelligence and autonomous driving, where high levels of safety, transparency, and interpretability are critical, while needing to handle highly complex environments.
Moreover, this modular characteristic facilitates human expert involvement in the policy discovery process. In DRL, the tight coupling of the entire network makes it challenging to modify specific control preferences. MLES, however, allows experts to inspect and intervene at any stage of the policy discovery process. If a deficiency is identified, the expert can modify the corresponding part of the policy and reintroduce it into the policy pool. This enables continuous expert-assisted refinement. A further advantage is that experts can simply identify deficiencies and suggest directions for improvement in a prompt, allowing MLES to autonomously implement targeted optimizations.
Easier knowledge transfer and reuse
A key advantage of MLES is that the discovered policies are encoded as executable code, which greatly simplifies knowledge transfer and reuse. In DRL, transferring knowledge often involves complex mechanisms like transfer learning or fine-tuning, which can be resource-intensive and difficult to implement. MLES bypasses these challenges by directly encoding prior knowledge into code. This approach not only makes knowledge transfer easier but also accelerates the policy deployment process. Additionally, LLMs can assist in this translation process, further reducing the time and effort required. As a result, MLES is particularly beneficial for tackling problems that involve well-established expert knowledge.
Moreover, this ease of transfer extends to addressing generalization challenges. In many tasks, different instances of the same problem may require subtle adjustments to the policy for optimal performance. Despite these variations, the core logic of the policy generally remains consistent. As such, instead of starting from scratch for each new task, MLES can adapt and fine-tune existing policies with minimal adjustments, ensuring continued high performance. A promising future direction for MLES is the collaborative evolution of policies for instances across multiple distributions, facilitating knowledge sharing of high-performing policies across different scenarios during the evolutionary process. This approach will be demonstrated in future work.
With the integration of visual feedback-driven behavior analysis, MLES mimics how human experts design policies—actively evaluating and adjusting policies in response to feedback. The key advantage is that MLES operates fully autonomously and can scale indefinitely, given sufficient computational resources. This scalability significantly accelerates the discovery of viable policies. In summary, MLES holds substantial potential to be a leading framework for interpretable policy discovery. We believe that MLES offers a new path forward for the field and invite the broader research community to explore and contribute to this exciting area.
A.2 Limitation
Construction of interpretable behavioral evidence
MLES relies on behavioral evidence to analyze policy behavior and guide targeted improvements. This requires the behavioral evidence to be sufficiently informative, capturing key performance aspects. In this paper, for the Lunar Lander and Car Racing tasks, we manually designed interpretable behavioral evidence from a human perspective and successfully applied it within MLES. However, for new tasks, human experts need to design task-specific evidence construction processes, introducing a certain amount of manual design work. To address this, we propose two potential solutions. First, adopting video-based behavioral evidence could provide rich, detailed information with minimal manual effort. However, this approach is computationally expensive and would require a reduction in MLLM costs before it becomes widely applicable. Second, extending MLES to support the automatic design of behavioral evidence. Prior to policy discovery, MLES could automatically design evidence construction pipelines, thereby generating evidence that facilitates the understanding of the task and the policy’s behavior by MLLMs, thus reducing the manual effort required.
Dependency on multimodal large language models.
The performance of MLES’s policy discovery process is inherently tied to the capabilities of the underlying MLLMs. Specifically, the analysis of behavioral evidence relies on the MLLMs’ visual understanding, while policy generation depends on their knowledge base and programming capabilities. As MLLMs continue to evolve, we anticipate that improvements in their capabilities will lead to enhanced policy discovery in MLES, allowing it to scale to more complex tasks and improve the overall effectiveness of the framework.
Higher cost compared to traditional deep reinforcement learning.
Another limitation of MLES is its higher computational cost due to the need to interact with MLLMs. While visual feedback-driven behavior analysis improves the policy discovery efficiency, it also increases the cost per policy generation. To mitigate this, we suggest expressing richer information within a single image (as demonstrated in this paper), thereby reducing the number of tokens used and lowering computational costs. Moreover, as MLLM technology advances and costs decrease, the overall cost of policy design within MLES will be significantly reduced.
Appendix B Details of Quantitative Metrics in MLES
B.1 Quantitative Metrics for Policy Evaluation in MLES
@@
This approach offers a practical advantage in that it eliminates the need for frequent manual design of intermediate rewards. We only need to focus on transforming the ultimate task goal into an evaluable metric, which can then support automated policy discovery. This not only simplifies the process but also alleviates the computational burden associated with reward design. However, this comes with the challenge of lacking direct evaluation of intermediate behaviors, which may potentially lead to issues such as reward hacking. To address this limitation, MLES incorporates interpretable behavioral evidence (IBE), which supplements the evaluation of intermediate actions. A more detailed discussion of this can be found in Section 3.3 and Appendix C.
@@
B.2 Quantitative Metrics for Benchmark Tasks
@@
Appendix C The Formats and Necessity of Interpretable Behavior Evidence
This section provides a detailed discussion of the potential formats of Interpretable Behavior Evidence (IBE) and the construction techniques involved, along with the specific forms adopted in this work and the rationale behind our choices.
C.1 Formats of Interpretable Behavior Evidence
The primary objective of IBE is to represent the behavior patterns associated with a given policy, thereby helping MLLMs in identifying policy shortcomings and facilitating subsequent improvements. IBE serves as a crucial supplement to the evaluation of intermediate actions, addressing the inherent limitations of purely quantitative metrics such as episode reward and success rate, which often fail to provide granular insights into policy execution.
The precise format of IBE is secondary to the quality and interpretability of the information it conveys, and how effectively this information can be leveraged by MLLMs. MLLM training paradigms, relying on human-provided images and corresponding natural language descriptions, progressively align model capabilities with human understanding. As evidenced by related work, current MLLMs demonstrate a level of visual comprehension closely mirroring human perception, enabling them to interpret information in a semantically meaningful way.
Figure 6: Examples of IBE in the Lunar Lander task. (a) Trajectory videos visualizing the agent’s motion over time. (b) Pose overlays summarizing sequential states in a single image. (c) State traces presenting a symbolic representation of the agent’s behavior.
Figure 6 illustrates three potential IBE formats within the context of the Lunar Lander task: videos, images, and textual state traces. While these formats convey the same underlying information, their interpretability and computational overhead for MLLM input vary significantly. Videos offer a complete visualization of the entire process but incur substantial computational expense. With appropriate preprocessing, images can effectively convey the control process while maintaining relatively low computational costs. Conversely, textual descriptions, though offering the lowest computational cost, are less intuitive for both human and MLLM interpretation, demanding greater cognitive effort for comprehension.
@@ C.1 Formats of Interpretable Behavior Evidence
Furthermore, real-world control tasks, such as bipedal robot walking, drone stabilization, or autonomous driving, frequently involve complex dynamics that are inherently difficult to articulate or fully capture through textual means. Visual formats are therefore superior for conveying these complex behaviors. In our experiments, to strike a balance between comprehensibility for both humans and MLLMs and computational efficiency, we opt to employ images as the primary format for IBE. The specific methodologies for constructing them are detailed in the subsequent section.
C.2 IBE for Two Benchmarks
This section details the construction of IBE for Lunar Lander and Car Racing tasks, addressing their unique characteristics.
@@ C.2 IBE for Two Benchmarks
The Lunar Lander task features a static background, with only the lander object in motion. Given this setup, we can effectively capture the entire behavioral pattern by using a frame stacking technique with transparency. As illustrated in Fig. 7, this process involves selecting frames at fixed intervals and applying a transparency effect to each one before stacking them. The resulting IBE is a single image that visualizes the lander’s pose and trajectory at key moments, effectively compressing the time-series information of an entire episode into a compact, visual representation.
Figure 7: Frame stacking for IBE construction for the Lunar Lander task
Figure 8: Trajectory mapping for IBE construction for the Car racing task
@@
The Car Racing task is a top-down environment where the camera continuously follows the car, and the background is dynamic. For this reason, the frame stacking method is not applicable. Instead, we propose an IBE construction method based on a global map and trajectory mapping, as shown in Fig. 8. The process begins by obtaining the complete coordinates of the track from the environment’s backend to construct a global track map. The car’s movement trajectory is then recorded throughout the episode and plotted onto this map upon completion. However, a simple trajectory alone is insufficient for explaining the policy’s decision-making process.
@@
C.3 Necessity of IBE
This section presents a qualitative analysis of IBE’s necessity, highlighting how it addresses a key limitation of quantitative metrics: their inability to evaluate the intermediate actions.
Fig. 9 illustrates the performance of policies designed by EoH and MLES on the same Car Racing track. Both policies achieve a perfect score, completing 100% of the track. However, a closer inspection of their trajectories reveals a critical difference. The EoH policy’s trajectory deviates significantly from the track on the final turn, only completing the course by executing a large corrective maneuver. This behavior is a classic example of ”reward hacking,” a common issue in RL where agents exploit flawed reward functions to achieve a high score through suboptimal or unintended behaviors (Skalse et al. 2022). In contrast, the MLES policy’s trajectory is both rational and efficient, closely mirroring a human driver’s racing line.
@@ C.3 Necessity of IBE
In our approach, MLES utilizes IBE as a crucial supplement to quantitative metrics, allowing MLLMs to perform in-depth behavioral analysis. This qualitative input enables the system to identify and proactively rectify these ”fake-excellent” policies. Furthermore, for policies that are sound but underperforming, IBE allows the MLLM to provide targeted feedback for specific improvements, thereby accelerating the policy discovery process. This highlights the indispensable role of IBE in the LES paradigm for achieving robust and generalizable policy learning.
Appendix D Additional experiment results
D.1 Hyperparameters for Experimental Setup
This subsection presents the hyperparameters used in our experiments involving MLES, EoH, DQN, and PPO. The experiments were implemented in Python and executed on a single CPU (Intel i9-13980HX) with 32GB of RAM. Table 3 details the hyperparameters for MLES and EoH, where identical settings were applied to ensure a fair comparison, based on the configurations from the EoH paper (2024a). Tables 4 and 5 present the hyperparameter settings for DQN and PPO across two tasks, respectively. Tables 6 and 7 show the network architectures used for DQN and PPO on two benchmarks. To ensure a fair comparison, we maintained identical network structures (with approximately equal parameter counts) for DQN and PPO in terms of feature processing and decision-making, excluding the output layers.
Hyperparameter Parameter Description Value
$m$ Number of parents selected in each evolution step 2
NN Population size 16
LLM Version of LLM used in the evolutionary operators GPT-4o-mini
Table 3: Hyperparameter settings for MLES and EoH.
Hyperparameter Lunar Lander Car Racing
Action Space Dimension 4 9 (Discrete actions)
Replay Buffer Capacity 10,000 30,000
Batch Size 128 128
Discount Factor 0.99 0.98
Initial $\epsilon$ 1.0 1.0
Minimum $\epsilon$ 0.01 0.1
$\epsilon$ Decay Rate 0.9995 0.995
Target Network Update Rate ($\tau$) 0.01 0.01
Optimizer Type Adam Adam
Learning Rate 0.001 0.0005
Table 4: DQN Hyperparameters for Lunar Lander and Car Racing Tasks
Parameter Lunar Lander Car Racing
Discount Factor 0.99 0.99
Generalized Advantage Estimation Parameter 0.95 0.95
PPO Clip 0.2 0.2
Entropy Coefficient 0.01 0.01
Critic Coefficient 0.5 0.5
Max Gradient Norm 0.5 0.5
Batch Size 256 256
Mini Batch Size 64 64
Learning Rate 3e-5 3e-5
Table 5: PPO Hyperparameters for Lunar Lander and Car Racing Tasks
Component Layer Type DQN Configuration PPO Configuration
Input Layer State Dimension 8 (environment observation space)
[…]
Output Layer Policy Head Linear(4)
Value Head N/A Linear(1)
Activation Hidden Layers ReLU
Output Layer Linear
Parameters Total Parameters  269,316  269,829
Table 6: Network Architectures for DQN and PPO on Lunar Lander Task
Component Specification DQN PPO
Layer 2 Conv2d(32→64, kernel=4×4, stride=2) → ReLU
Layer 3 Conv2d(64→64, kernel=3×3, stride=1) → ReLU
Value Head Architecture N/A Linear(512→256) → ReLU → Linear(256→1)
Activation Hidden ReLU
Output Linear Policy: Tanh/Sigmoid, Value: Linear
Parameters Total 4,277,417 4,536,484
Table 7: Network Architectures for DQN and PPO on Car Racing Task
D.2 Statistical Analysis of Comparative Evaluations
Table 8 presents the average, standard error of the mean (SEM), and best results across five runs on training instances, complementing the data shown in Table 1. With the aid of visual feedback-driven behavior analysis, MLES demonstrates superior stability and convergence compared to EoH. In the Car Racing task, both MLES and PPO are able to design policies that achieved a perfect score across all 10 test instances. Notably, as shown in Table 7, while PPO utilizes a significantly larger number of parameters, the policies designed by MLES involve considerably less computational complexity, making them more suitable for deployment in real-world applications.
Method Lunar Lander Car Racing
Ave SEM Best Ave SEM Best
DQN 1.017 $\pm$0.022 1.066 79.777 $\pm$2.635 91.182
PPO 1.032 $\pm$0.026 1.076 99.212 $\pm$0.390 100.000
EoH 1.053 $\pm$0.021 1.085 89.808 $\pm$1.553 96.675
MLES 1.090 $\pm$0.005 1.098 98.704 $\pm$0.719 100.000
Table 8: Average, best results, and SEM from five runs on training instances
D.3 Influence of Behavioral Evidence and Prompt Engineering on Policy Discovery
This section investigates the influence of different forms of IBE and prompt engineering strategies on the policy discovery process. We employ M1 as the test operator and construct various few-shot prompts, incorporating diverse forms of IBE and specific instructions. Specifically, we evaluate the following five configurations:
To quantitatively assess the impact of these different IBE forms and prompt instructions, we initialized our experiments with five distinct populations, each exhibiting significant evolutionary potential. These populations are carefully selected from an established policy evolution process for the Lunar Lander task by MLES. Specifically, we identify five generations that have previously demonstrated substantial progress and select their immediate preceding populations as our initial populations. This selection criterion ensures that our starting points possess considerable room for improvement, allowing for clearer observation of the operators’ effects. Each M1 operator variant is applied to these populations for two generations. The entire experiment is repeated five times to ensure statistical robustness. We measure the impact by observing both the average performance improvement of the population and the improvement of the best policy within each population.
M1 M1_M M1_M M1_T M1_M
Population improvement +19.43% +24.55% +20.21% +20.29% +19.73%
Best policy improvement +1.18% +4.39% +2.37% +3.64% +2.95%
1. Integrating IBE significantly boosts policy discovery efficiency.
A foundational finding is that the incorporation of IBE consistently leads to improved performance compared to the baseline M1 operator. The baseline M1, without any IBE, achieves a population improvement of +19.43% and a best policy improvement of +1.18%. In contrast, all configurations that utilize IBE (M1_M, M1_M, M1_T, M1_M) exhibit higher gains in both population and best policy performance. For instance, M1_M demonstrates the highest population improvement at +24.55% and also the most significant best policy improvement at +4.39%. This clearly indicates that providing MLLMs with relevant behavioral evidence significantly aids in fine-grained policy evaluation and enables more targeted policy optimization. The evidence acts as a crucial guide, steering the MLLM toward more effective policy spaces.
2. The impact of interpretability and richness of IBE.
Both M1_M and M1_T demonstrate substantial improvements in best policy performance, with +4.39% and +3.64%, respectively. While both forms of IBE prove beneficial, the visual representation of the lander’s behavior and posture in images is arguably more expressive, intuitive, and easier for the MLLM to interpret than raw textual state traces. This superior interpretability and rich expressiveness enable the MLLM to analyze policy behavioral patterns more deeply. These results support our claim, as discussed in Appendix C.1, that the most crucial aspect of IBE is the quality and inherent interpretability of the information it conveys, along with how effectively this information can be leveraged by MLLMs for comprehensive behavioral analysis and subsequent policy refinement. The richer, more direct semantics embedded in visual data appear to offer a distinct advantage.
3. The benefit of direct image input.
Comparing M1_M (direct image input) with M1_M (two-stage processing where an image is first described by an MLLM into text, then passed to the policy generator), we observe a clear performance disparity. M1_M achieves a population improvement of +24.55% and a best policy improvement of +4.39%, whereas M1_M yields slightly lower gains at +19.73% for population and +2.95% for best policy. When an image transforms into a textual description by another model, even if that description is ”detailed,” there is an inherent risk of information loss or misinterpretation. For an MLLM serving as a policy generator, directly accessing the raw image appears more advantageous than relying on a cascaded interpretation. This suggests that the rich, nuanced information present in the original image may be difficult to fully capture and convey through an intermediate textual representation, at least with current LLM capabilities. This outcome highlights the value of direct multimodal processing for inherently multimodal tasks, as it avoids potential bottlenecks and accumulated errors introduced by an intermediate textualization step.
4. The advantage of explicit instructions for behavioral analysis.
The contrast between M1_M and M1_M strongly highlights the advantage of providing explicit instructions for behavior analysis within the prompt. M1_M consistently performs better in both population (+24.55%) and best policy (+4.39%) improvement compared to M1_M (+20.21% and +2.37% respectively). Both configurations receive image input, but M1_M explicitly instructs the LLM to provide a detailed description and analysis of the image, while M1_M does not. This gap emphasizes that while MLLMs can implicitly derive information from raw image input to aid inference, explicit directives to analyze IBE can provide rich, relevant context. This promotes the activation of pertinent knowledge within the MLLM. By encouraging a ”thinking aloud” process, such instructions enable the LLM to synthesize more comprehensive and actionable insights, which are subsequently utilized to generate more effective policy modifications. This finding is consistent with observations in Chain of Thought and similar works (Yu et al. 2025; Wang and Venkatesh 2025).
In summary, these experiments underscore that both the form of IBE and the strategy of prompt engineering are crucial factors in maximizing the effectiveness of MLES-driven policy discovery. A well-crafted context, combining expressive and interpretable IBE with clearly instructed analytical processes during MLLM inference, significantly improves reasoning and enables more targeted and effective policy optimization. Our findings suggest that for complex tasks requiring a nuanced understanding of behavior, direct multimodal input, coupled with explicit analytical prompts, represents a superior approach.

### v2

@@ B.5 Generalization via Policy Ensemble
The results are highly encouraging. Despite the simplicity of the ensemble method, it produced a significant improvement in test performance for both MLES and EoH. It is worth clarifying that a slight decline in training performance is also observed after ensembling. This outcome is a well-known characteristic of the bias-variance tradeoff, where ensembling acts as a form of regularization to improve generalization, sometimes at the cost of slightly increased bias on the training set.
Crucially, the complementarity of the policies in the population is achieved organically, without employing any explicit diversity-preservation techniques common in the EC field. This indicates a substantial, yet-to-be-tapped potential for further improving MLES to address generalization issues. Looking forward, we believe this provides a promising direction for future work, which could focus on explicitly cultivating a “specialist team” of policies. This might involve exploring more advanced techniques, such as:
  • Feature-Aware Instance Clustering: Automatically grouping similar problem instances based on their features to train or select specialized policies for each cluster.
  • Cooperative Co-evolution: Designing efficient evolutionary mechanisms that encourage the policy population to collectively cover the full spectrum of training instances, rewarding policies for succeeding where others fail.
  • Adaptive Collaborative Decision-Making: Creating dynamic ensembling mechanisms where policies “vote” on the final action with varying confidence levels based on the current state.
Overall, while not the primary focus of this paper, MLES demonstrates inherent properties and substantial potential that could be harnessed to tackle the challenge of generalization.
Appendix C Details of Quantitative Metrics in MLES
C.1 Quantitative Metrics for Policy Evaluation in MLES
@@
This approach offers a practical advantage in that it eliminates the need for frequent manual design of intermediate rewards. We only need to focus on transforming the ultimate task goal into an evaluable metric, which can then support automated policy discovery. This not only simplifies the process but also alleviates the computational burden associated with reward design. However, this comes with the challenge of lacking direct evaluation of intermediate behaviors, which may potentially lead to issues such as reward hacking. To address this limitation, MLES incorporates behavioral evidence (BE), which supplements the evaluation of intermediate actions. A more detailed discussion of this can be found in Section 2.3 and Appendix D.
@@
C.2 Quantitative Metrics for Benchmark Tasks
@@
Appendix D The Formats and Necessity of Behavior Evidence
This section provides a detailed discussion of the potential formats of Behavior Evidence (BE) and the construction techniques involved, along with the specific forms adopted in this work and the rationale behind our choices.
D.1 Formats of Behavior Evidence
The primary objective of BE is to represent the behavior patterns associated with a given policy, thereby helping MLLMs identify policy shortcomings and facilitate subsequent improvements. BE serves as a crucial supplement to the evaluation of intermediate actions, addressing the inherent limitations of purely quantitative metrics such as episode reward and success rate, which often fail to provide granular insights into policy execution.
The precise format of BE is secondary to the quality and interpretability of the information it conveys, and how effectively this information can be leveraged by MLLMs. MLLM training paradigms, relying on human-provided images and corresponding natural language descriptions, progressively align model capabilities with human understanding. As evidenced by related work, current MLLMs demonstrate a level of visual comprehension closely mirroring human perception, enabling them to interpret information in a semantically meaningful way.
Figure 7: Examples of BE in the Lunar Lander task. (a) Trajectory videos visualizing the agent’s motion over time. (b) Pose overlays summarizing sequential states in a single image. (c) State traces presenting a symbolic representation of the agent’s behavior.
Fig. 7 illustrates three potential BE formats within the context of the Lunar Lander task: videos, images, and textual state traces. While these formats convey the same underlying information, their interpretability and computational overhead for MLLM input vary significantly. Videos offer a complete visualization of the entire process but incur substantial computational expense. With appropriate preprocessing, images can effectively convey the control process while maintaining relatively low computational costs. Conversely, textual descriptions, though offering the lowest computational cost, are less intuitive for both human and MLLM interpretation, demanding greater cognitive effort for comprehension.
@@ D.1 Formats of Behavior Evidence
Furthermore, real-world control tasks, such as bipedal robot walking, drone stabilization, or autonomous driving, frequently involve complex dynamics that are inherently difficult to articulate or fully capture through textual means. Visual formats are therefore superior for conveying these complex behaviors. In our experiments, to strike a balance between comprehensibility for both humans and MLLMs and computational efficiency, we opt to employ images as the primary format for BE. The specific methodologies for constructing them are detailed in the subsequent section.
D.2 BE for Two Benchmarks
This section details the construction of BE for Lunar Lander and Car Racing tasks, addressing their unique characteristics.
@@ D.2 BE for Two Benchmarks
The Lunar Lander task features a static background, with only the lander object in motion. Given this setup, we can effectively capture the entire behavioral pattern by using a frame stacking technique with transparency. As illustrated in Fig. 8, this process involves selecting frames at fixed intervals and applying a transparency effect to each one before stacking them. The resulting BE is a single image that visualizes the lander’s pose and trajectory at key moments, effectively compressing the time-series information of an entire episode into a compact, visual representation.
Figure 8: Frame stacking for BE construction for the Lunar Lander task
Figure 9: Trajectory mapping for BE construction for the Car racing task
@@
The Car Racing task is a top-down environment where the camera continuously follows the car, and the background is dynamic. For this reason, the frame stacking method is not applicable. Instead, we propose a BE construction method based on a global map and trajectory mapping, as shown in Fig. 9. The process begins by obtaining the complete coordinates of the track from the environment’s backend to construct a global track map. The car’s movement trajectory is then recorded throughout the episode and plotted onto this map upon completion. However, a simple trajectory alone is insufficient for explaining the policy’s decision-making process.
@@
D.3 Necessity of BE
This section presents a qualitative analysis of BE’s necessity, highlighting how it addresses a key limitation of quantitative metrics: their inability to evaluate the intermediate actions.
@@ D.3 Necessity of BE
Fig. 10 illustrates the performance of policies designed by EoH and MLES on the same Car Racing track. Both policies achieve a perfect score, completing 100% of the track. However, a closer inspection of their trajectories reveals a critical difference. The EoH policy’s trajectory deviates significantly from the track on the final turn, only completing the course by executing a large corrective maneuver. This behavior is a classic example of “reward hacking,” a common issue in RL where agents exploit flawed reward functions to achieve a high score through suboptimal or unintended behaviors Skalse et al. (2022). In contrast, the MLES policy’s trajectory is both rational and efficient, closely mirroring a human driver’s racing line.
@@
In our approach, MLES utilizes BE as a crucial supplement to quantitative metrics, allowing MLLMs to perform in-depth behavioral analysis. This qualitative input enables the system to identify and proactively rectify these “fake-excellent” policies. Furthermore, for policies that are sound but underperforming, BE allows the MLLM to provide targeted feedback for specific improvements, thereby accelerating the policy discovery process. This highlights the indispensable role of BE in the LES paradigm for achieving robust and generalizable policy learning.